The product of two consecutive integers is 5 more than their sum. Find the integers.
Answer by mananth(16946)

let x & x+1 be the numbers
x(x+1) = x+x+1+5
x^2+x=2x+6
x^2-x-6=0
x^2-3x+2x-6=0
x(x-3)+2(x-3)=0
(x+2)(x-3)=0
x=-2 OR 3
The numbers are
3& 4 OR -2,-1